OpenStudy (anonymous):

1. A model is made of a car. The car is 7 feet long, and the model is 16 inches long. What is the ratio of the length of the car to the length of the model? (1 point) 4:21 7:16 16:7 21:4 2. The measure of two complementary angles is in the ratio 1:9. What are the degree measures of the two angles? (1 point) 9° and 81° 10° and 80° 18° and 162° 20° and 160° 3. A salsa recipe uses green pepper, onion, and tomato in the extended ratio 2:5:9. How many cups of onion are needed to make 96 cups of salsa? (1 point) 12 5 30 54

OpenStudy (anonymous):

7 feet : 16 inches 7x12 inches : 16 inches 84 inches : 16 inches 84:16 divide both sides by 4 21 : 4

OpenStudy (anonymous):

ohh thank u can u help me witth the rest

OpenStudy (anonymous):

x + 9x = 90 degrees 10x = 90 x = 9 x is 9 .... 9x is 9*9 or 81 9 and 81

OpenStudy (anonymous):

2+5+9=16 96 / 16 = 6 5 x 6 = 30

OpenStudy (anonymous):

4. What is the solution to the proportion? five over eight equals m over 40 (1 point) 1 over 25 25 5 1 5. Given the proportion a over b= 7 over 13, which ratio completes the equivalent proportion a over 7= ? (1 point) a over 13 b over 13 7 over 13 13 over b

OpenStudy (anonymous):

You're gonna have to do SOME of the work here 5/8 = m/40 First step, multiply both sides by 40

OpenStudy (anonymous):

200 over 330 =m

OpenStudy (anonymous):

help pleas

OpenStudy (anonymous):